# Double white dwarf binary population in MOCCA star clusters

### v2

@@ 3.3 Star cluster and isolated binary evolution
For each initial setup, we ran two simulations; one with dynamical interactions between objects and one run with no dynamical interactions, i.e. pure binary/stellar evolution where all objects in our initial star cluster model are evolved in isolation. This allows us to understand the impact of dynamics on the primordial binary population and DWD formation. In addition, whenever an object is ejected from a cluster, we used bse to continue the evolution of this object. Furthermore, we split up primordial and dynamically formed binaries in the cluster simulation with dynamics. This gives the following 4 groups:
  • Cluster-primordial: DWDs with progenitors born together found inside an evolving cluster.
  • Cluster-dynamical: DWDs that were formed in a dynamical interaction found inside an evolving cluster.
  • Standalone evolution: DWDs evolved in isolation
  • Escapers: escaping DWDs from the cluster simulations. Evolved as isolated binaries after escape.
@@ 3.4 Further calculations
There is evidence for natal kicks on WDs (Davis et al. 2008; Hamers & Thompson 2019) that could lead to the destruction of fragile wide binaries. The magnitude of these kicks are believed to be on the order of $\sim 0.75$ km/s (El-Badry & Rix 2018). This is not included in our simulations and thus we apply the energy related to a kick of $0.75$ km/s for each star to the binding energy of the binary. We check if this causes the binary to dissolve, if it does, we remove this binary from the sample. If not, we calculate the new semi-major axis related to the new binding energy. To calculate the absolute magnitude of the WDs from mocca we use the FSPS [^2] https://github.com/cconroy20/fsps code (Conroy et al. 2009; Conroy & Gunn 2010). For one binary, this code can calculate the absolute magnitudes of both components in a binary using the mass, radius, luminosity and effective temperature. This provides us with absolute magnitudes for different instruments and filters, such as the Gaia G band.
@@ 3.5 Observational selection effects
Binary stars can be detected through different methods depending on the distance between the two components. Close binaries are detected through e.g. photometric surveys, spectroscopic observations or X-ray observations while wide binaries are too far apart to reliably use photometric data. Instead, if two wide separated stars have a common proper motion across the sky or similar radial velocities they can be classified as a binary candidate and in more detailed follow-up observations this can be confirmed. Due to the different methods of obtaining observational data, there are also different limitations. We need to consider these observational selection effects when comparing our data. Thus we apply different filtering criteria on the close and wide binaries respectively. These criteria are explained in the following sections.
@@ 3.5.1 Close binaries
  • P ¡ 1.5 day
  • G ¡ 11.24 mag where G is the combined absolute magnitude of the two WDs:
    $G=-2.5\log(10^{-0.4G_{1}}+10^{-0.4G_{2}})$ (2)
    where $G_{1}$ and $G_{2}$ is the absolute magnitudes in the G-band of the primary and secondary WD.
@@ 3.5.2 Wide binaries
  • Semi-major axis: $a>100$ AU
  • Absolute magnitude: $G<16.1$ mag
  • Addition of kinetic energy corresponding to kicks of WDs to the binding energy of the binary (see Sect. 3.2).
@@ 3.6 Projection
It is often impossible to determine the semi-major axis of wide binaries from observational data. Instead, the projected separation between the two binary components is the parameter that can be determined. Due to this, to compare the binaries from our simulations to the observed binaries, we need to project our semi-major axis. The projection method we use is derived from chapter 2 of Hilditch (2001).
Assuming that we put one of the binary components at $(x,y,z)=(0,0,0)$. And assume that the xy plane is perpendicular to the line of sight of the observer, thus we cannot see the z-axis. We can calculate the projected position onto the x and y axes:
@@
$r\sin\theta[-\sin\Omega\sin\omega+\cos\Omega\cos\omega\cos i]$
where $\Omega$ is the longitude of the ascending node, $\omega$ is the longitude of periastron and $i$ is the inclination of the orbit (see Fig. 2.5 in Hilditch (2001)). At pericenter and apocenter ($\theta=0$ and $\theta=\pi$ respectively) We have $\sin(0)=\sin(\pi)=0$:
$x=r\cos\theta[\cos\Omega\cos\omega-\sin\Omega\sin\omega\cos i]$ (4)
$y=r\cos\theta[\sin\Omega\cos\omega+\cos\Omega\sin\omega\cos i]$
To get an average value of the projection fraction $f_{p}$, the average value that the semi-major axis is changed when projected onto the sky, we integrate over all angles; $\Omega$ from 0 to $\pi$, and $\omega$ $i$ from 0 to $\pi/2$. This gives
$x=y\approx-0.2584\,r$ (5)
The projection fraction is thus
$f_{p}=\sqrt{x^{2}+y^{2}}=0.842$ (6)
which is used to calculate the projected distance as:
$a_{\rm proj}=a\cdot f_{p}$ (7)
where a is the semi-major axis of the binary.
@@ 4 Results
We start the presentation of our results by comparing the projected separation (calculated according to Sect. 3.6) with the mass of the WDs. For the observed close binaries we only consider the mass of the lower mass WD since the higher mass WD is typically not visible and the errors in the masses are usually very large. For all other data sets, both observations and simulations, we consider the mass of both WDs. Since observed close binaries cannot be resolved individually, there is no observational data for the separation of these binaries. Instead, we assume that the eccentricity of these binaries is 0 due to circularization and calculate the semi-major axis using Kepler’s third law.
@@
In Fig. 1 we show the projected separation plotted against the mass of the WDs for clusters with 10% and 95% respectively initial binary fraction at 9 Gyrs snapshot. This plot shows the whole range of separations, from close to wide, for the four different groups defined in Sec. 3.3.
Due to the many similarities and lack of clear differences in the overall picture for 10% and 95% binary fraction we will discuss both datasets together. We do not filter on magnitude in this section since the magnitude limits are different in different regions of the plot.
@@
A noticeable feature of the in-cluster binaries is a gap in separations between 0.6 AU and 25 AU. This gap only exists for in-cluster binaries while for both escapers and isolated ones many systems are occupying this region in the parameter space. The reason for this gap is dynamics, more specifically, the threshold between soft and hard binaries lies in this gap. Binaries below this threshold will, on average, become harder due to dynamical interactions while binaries above will, on average, become softer as a result of dynamical interactions. This pushes binaries away from the threshold and creates this gap. This is one example of how important dynamical interactions can be for binaries in these dense environments.
@@
The distribution of escapers is very similar to isolated binaries on a statistical scale. However, a very noticeable trend is that for all times and across all GC models we have far more escaped DWDs than isolated DWDs. The exact number depends on initial conditions and evolution time but we found that usually this ratio is between 1 and 3. This seems to indicate that dynamics play an important role in the formation of these binaries and without dynamics the formation of these binaries are not as effective as when dynamical interactions are included.
@@
There is also a big difference in the number of escapers and in-cluster binaries. The reason for this is most likely due to stellar evolution leading to rapid mass loss from the evolution of massive stars in the first hundred Myr of the cluster evolution. When the stars are initially evolving and losing mass the tidal radius is decreased strongly and rapidly. This means that many binaries are suddenly outside of the tidal radius and thus counted as escapers. The binaries can later evolve into DWDs. We discuss this in more detail in Sect. 3.2.
@@
For very low separations, there is a hook shaped formation, particularly for escapers and in-cluster binaries. This is related to cataclysmic variables that have later evolved into a DWD system. The fact that we have far fewer systems in this formation for isolated binaries once again shows the importance of dynamical interactions leading to their formation. For escapers in the 10% initial binary fraction datasets there are also a few lines going towards lower separations. These are also related to cataclysmic variables.
@@
The observations of DWDs are currently limited to either spectroscopic binaries or wide binaries DWDs. While themocca data spans a wide range of separations and masses while observations are either for very close binaries or for very wide. We can clearly see that the observations are only including a very small part of the whole population. Because of this, we cannot do a direct comparison between mocca data and observations. We split our results into close and wide binaries and restrict our mocca data as explained in sections 3.5.1 and 3.5.2. The results for close binaries are found in Sect. 4.1 and for wide binaries in Sect. 4.2.
@@
Figure 1: Projected separation plotted against mass of the lower mass WD in the binary for 10% (panel a) and 95% (panel b) initial binary fraction for the whole parameter space. . The numbers inside the square brackets in the legends show the total number of WD binaries in the runs. The different colours correspond to the different datasets: orange for escapers, blue for standalone binaries, red for in-cluster primordial binaries, green for in-cluster dynamically formed binaries, dark red for observations of close binaries and purple for observations of wide binaries. In addition to this, we have four different types of markers: stars for observed WDs, circles for helium WDs, diamonds for carbon oxygen WDs, pluses for oxygen neon WDs
@@ 4.1 Close binaries
In order to go into more detail when looking at the close binaries we isolate binaries according to the criteria described in 3.5.1 and will thus only focus on a small subset of the total population. We plot the projected separation against mass of the lower mass WD in Fig. 2 and 3. We only include the mass of the lower mass WD since, due to the observational technique, the errors in mass of the higher mass WD are often very large. Due to our limit on magnitude, we see that over time, higher mass WDs are not visible due to the cooling process of WDs. Already at 5 Gyr, all WDs with masses higher than $\sim 0.55$M are now undetectable.
4.1.1 10% initial binary fraction
@@
We will start by looking at the 10% initial binary fraction plots in Fig. 2. At 2 Gyr the separations from all mocca datasets are slightly too wide and the masses too high. At 5 Gyr, escapers and isolated binaries agree very well with observations in separation. The in-cluster binaries at this time are biased towards lower separations than the other data-sets. At 9 Gyr, not much changes compared to 5 Gyr, the maximum mass is slightly lower due to more binaries being non-detectable due to increasing magnitude as they cool down. However, for all snapshots, the masses are not well reproduced. We are not able to produce very low mass WDs with mass below 0.2 M in our simulations. Instead, our close WDs usually have a mass between 0.25 and 0.4 M. There are a few observed WDs in this mass range as well but the large majority is smaller mass.
@@
These ELMs cannot form through common envelope evolution which produces a problem for our comparisons. Currently, bse does not handle this kind of evolution well. There is a large uncertainty in the magnetic braking, mass stability criteria and mass transfer modelling. More precisely, to form these objects we need dynamically stable mass transfer (e.g. Belloni & Schreiber 2023b). During this stable mass transfer involving a Roche-lobe filling subgiant or an unevolved red giant star, due to strong magnetic braking, the evolution is convergent and the binary evolves towards shorter orbital periods. This kind of mass transfer is not handled by bse or any other population synthesis code at this moment in time. In order to be able to properly reproduce the distribution of mass for these kinds of ELMs, upgrades to our bse are needed, however, that is outside of the scope of this paper.
@@
Escaping binaries are able to form lower mass WDs than in-cluster or isolated binaries which can be seen as a sub-population between 0.3 and 0.45 M. This is most likely due to dynamical interactions before they escaped their host cluster. The in-cluster binaries that would form these low mass WDs are either escaping or dynamical interactions are causing an exchange, a break-up or a merger.
@@
At $\sim 0.51$ $\rm M_{\odot}$ we can see a soft limit for WD masses. There are WDs with lower mass than this but they are few. This is a limit in bse and a lower mass would require processes that are neither in bse or solved in theory. This includes, as mentioned above, uncertainty in magnetic braking, mass stability criteria and mass transfer modelling.
4.1.2 95% initial binary fraction
@@
For 95% initial binary fraction (Fig. 3) we have a similar scenario but there are small differences. At 2 Gyr we have a larger shift towards higher separations compared to 10% binary fraction. As expected, similarly to 10% binary fraction, we are not able to form the ELMs WDs that have been observed. At 5 and 9 Gyr, we have good agreement in separation between escapers, isolated binaries and observations. In-cluster binaries are biased towards lower separations.
@@
Looking at the number of close binaries we can see that the number of binaries is increasing with time for escapers and isolated binaries since over time more binaries are escaping their host cluster or the components of the binary are evolving into WDs. However, for in-cluster binaries, the number is decreasing due to escaping binaries and dynamical interactions causing break-ups, mergers or exchanges.
@@
An interesting feature is a triangular shaped gap in Fig. 3, panel (a). This gap occurs between 0.02 and 0.07 AU in projected separation and 0.8 and 1.1 $\rm M_{\odot}$ in mass. This gap is present for all datasets and the reason for it is currently unknown, although it is most likely related to the correlated distributions of the initial binary populations.
@@
Figure 2: Projected separation plotted against mass of the lower mass WD in the binary for close binaries inside clusters with 10% initial binary fraction. See Fig. 1 for an explanation of panels, markers and colours.
@@
Figure 3: Projected separation plotted against mass of the lower mass WD in the binary for close binaries inside clusters with 95% initial binary fraction. See Fig. 1 for an explanation of panels, colours and markers

### v3

@@ 3.3 Star cluster and isolated binary evolution
For each initial setup, we ran two simulations, one with dynamical interactions between objects and one run with no dynamical interactions (i.e., pure binary or stellar evolution where all objects in our initial star cluster model are evolved in isolation). This allows us to understand the impact of dynamics on the primordial binary population and DWD formation. In addition, whenever an object is ejected from a cluster, we used BSE to continue the evolution of this object. Furthermore, we split up primordial and dynamically formed binaries in the cluster simulation with dynamics. This gives the following four groups:
  • cluster-primordial: DWDs with progenitors born together found inside an evolving cluster;
  • cluster-dynamical: DWDs that were formed in a dynamical interaction found inside an evolving cluster;
  • standalone evolution: DWDs evolved in isolation;
  • escapers: DWDs that escaped from the cluster simulations, and that evolved as isolated binaries after escape.
@@ 3.4 Further calculations
There is evidence for natal kicks on WDs (Davis et al. 2008; Hamers & Thompson 2019) that could lead to the destruction of fragile wide binaries. The magnitude of these kicks is believed to be on the order of $\sim 0.75$ km/s (El-Badry & Rix 2018). This is not included in our simulations, and thus we add the energy related to a kick of $0.75$ km/s for each star to the binding energy of the binary. This can be seen as an upper limit scenario. We check whether this causes the binary to dissolve. If it does, we remove this binary from the sample. If not, we calculate the new semi-major axis related to the new binding energy.
To calculate the absolute magnitude of the WDs from MOCCA we use the FSPS [^2] https://github.com/cconroy20/fsps code (Conroy et al. 2009; Conroy & Gunn 2010). For one binary, this code can calculate the absolute magnitudes of both components in a binary using the mass, radius, luminosity, and effective temperature. This provides us with absolute magnitudes for different instruments and filters, such as the Gaia G band.
@@ 3.5 Observational selection effects
Binary stars can be detected through different methods depending on the distance between the two components. Close binaries are detected, for example, through photometric surveys, spectroscopic observations, or X-ray observations, while wide binaries are too far apart for us to reliably use photometric data. Instead, if two wide separated stars have a common proper motion across the sky or similar radial velocities, they can be classified as a binary candidate, and in more detailed follow-up observations this can be confirmed. Due to the different methods of obtaining observational data, there are also different limitations. We need to consider these observational selection effects when comparing our data. Thus, we apply different filtering criteria on the close and wide binaries. These criteria are explained in the following sections.
@@ 3.5.1 Close binaries
  • P ¡ 1.5 day;
  • $G_{\mathbf{abs}}$ 11.24 mag, where $G_{\mathbf{abs}}$ is the combined absolute magnitude of the two WDs:
    $G_{\mathbf{abs}}=-2.5\log(10^{-0.4G_{\mathbf{abs},1}}+10^{-0.4G_{\mathbf{abs},2}}),$ (2)
    where $G_{\mathbf{abs},1}$ and $G_{\mathbf{abs,}2}$ are the absolute magnitudes in the G band of the primary and secondary WD.
@@ 3.5.2 Wide binaries
  • semi-major axis: $a>100$ AU;
  • absolute magnitude: $G_{\mathbf{abs}}<16.1$ mag;
  • addition of kinetic energy corresponding to kicks of WDs to the binding energy of the binary (see Sect. 3.4).
@@ 3.6 Projection
It is often impossible to determine the semi-major axis of wide binaries from observational data. Instead, the projected separation between the two binary components is the parameter that can be determined. This means that to compare the binaries from our simulations to the observed binaries, we need to project our semi-major axis. The projection method we use is derived from chapter 2 of Hilditch (2001). The radius vector (r) joining the two binary components at a given time is at an angular position $\theta$, the true anomaly relative to the position of periastron of the orbit. This radius vector can be projected onto the x and y coordinate axes as
@@
$r\sin\theta[-\sin\Omega\sin\omega+\cos\Omega\cos\omega\cos i],$
where $\Omega$ is the longitude of the ascending node, $\omega$ is the longitude of periastron, and $i$ is the inclination of the orbit (see Fig. 2.5 in Hilditch 2001). Knowing the masses, semi-major axis, and eccentricity of the binary system, we can randomly sample the Keplerian orbital elements in Eq. 3 and the mean anomaly in order to use Kepler’s equation to find the corresponding eccentric anomaly and then the true anomaly. This allows us to obtain the Cartesian position coordinates of the two binary components. It is assumed that the observers line of sight is along the z direction, and thus with the position of the binary components we calculate the projected separation as
$\mathbf{a_{proj}=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}},$ (4)
where $x_{n}$ and $y_{n}$ are the x and y positions of the two respective binary components. To obtain an average projected separation, we do this 10000 times for each binary and use the average separation (for more details about this procedure, see Askar et al. 2018).
@@ 4 Results
We start the presentation of our results by comparing the projected separation (calculated according to Sect. 3.6) with the mass of the WDs. For the observed close binaries we only consider the mass of the lower mass WD since the higher mass WD is typically not visible and the errors on the masses are usually very large. For all other datasets, both observations and simulations, we consider the mass of both WDs. Since observed close binaries cannot be resolved individually, there is no observational data for the separation of these binaries; however, we know the period. Using the periods, we assume that the eccentricity of these binaries is 0 due to circularization, and calculate the semi-major axis using Kepler’s third law.
@@
In Fig. 1 we show the projected separation plotted against the mass of the WDs for clusters with a 10% and 95% initial binary fraction at 9 Gyrs snapshot. This plot shows the whole range of separations, from close to wide, for the four different groups defined in Sect. 3.3.
The differences in the overall picture are very small between a 10% and 95% binary fraction, and thus we discuss them together unless stated otherwise. Our magnitude limits are based on observational limitations, and thus we can only apply these limitations to separations where DWDs have been observed. In addition, the magnitude limits are different for close and wide binaries, which means that we are unable to filter based on magnitude for these two plots. In later sections we present the 10% and 95% binary fraction separately and we include the magnitude limits discussed in Sect. 3.5.
@@
A noticeable feature of the in-cluster binaries is a gap in separations between 0.6 AU and 25 AU. This gap only exists for in-cluster binaries, while for escapers and isolated binaries many systems occupy this region in the parameter space. The reason for this gap is dynamics; more specifically, the threshold between soft and hard binaries lies in this gap. Binaries below this threshold will, on average, become harder due to dynamical interactions while binaries above will, on average, become softer as a result of dynamical interactions. This pushes binaries away from the threshold and creates this gap. This is one example of how important dynamical interactions can be for binaries in these dense environments.
@@
The distribution of escapers is very similar to that of isolated binaries on a statistical scale. However, a very noticeable trend is that for all times and across all GC models we have far more escaped DWDs than isolated DWDs. The exact number depends on the initial conditions and the evolution time, but we found that usually the ratio is between 1 and 3. This seems to indicate that dynamics play an important role in the formation of these binaries, and without dynamics the formation of these binaries is not as effective as when dynamical interactions are included.
@@
There is also a big difference in the number of escapers and in-cluster binaries. The reason for this is most likely due to stellar evolution leading to rapid mass loss from the evolution of massive stars in the first 100 Myr of the cluster evolution. When the stars are initially evolving and losing mass the tidal radius decreases strongly and rapidly. This means that many binaries are suddenly outside of the tidal radius, and are thus counted as escapers. The binaries can later evolve into DWDs. We discuss this in more detail in Sect. 3.2.
@@
For very small separations, there is a hook-shaped formation, particularly for escapers and in-cluster binaries. This is related to cataclysmic variables that have later evolved into a DWD system. The fact that we have far fewer systems in this formation for isolated binaries once again shows the importance of dynamical interactions leading to their formation. For escapers in the 10% initial binary fraction datasets, there are also a few lines going toward smaller separations. These are also related to cataclysmic variables.
@@
The observations of DWDs are currently limited to either spectroscopic binaries or wide binary DWDs. While the MOCCA data spans a wide range of separations and masses, the observations are either for very close or for very wide binaries. We can clearly see that the observations are only including a very small part of the whole population. Because of this, we cannot make a direct comparison between the MOCCA data and observations. We split our results into close and wide binaries and restrict our MOCCA data, as explained in Sects. 3.5.1 and 3.5.2. The results for close binaries are found in Sect. 4.1 and for wide binaries in Sect. 4.2.
Refer to captionRefer to caption
@@
Figure 1: Projected separation plotted against mass of the lower mass WD in the binary for 10% (panel a) and 95% (panel b) initial binary fraction for the whole parameter space. The numbers inside the square brackets in the legends show the total number of WD binaries in the runs. The different colors correspond to the different datasets: orange for escapers, blue for standalone binaries, red for in-cluster primordial binaries, green for in-cluster dynamically formed binaries, dark red for observations of close binaries, and purple for observations of wide binaries. In addition, we have four different types of markers: stars for observed WDs, circles for helium WDs, diamonds for carbon oxygen WDs, pluses for oxygen neon WDs.
@@ 4.1 Close binaries
In order to go into more detail when looking at the close binaries, we isolate binaries according to the criteria described in 3.5.1, and thus only focus on a small subset of the total population. We plot the projected separation against mass of the lower mass WD in Figs. 2 and 3. We only include the mass of the lower mass WD since, due to the observational technique, the errors on the mass of the higher mass WD are often very large. Our limit on magnitude means that over time we see that higher mass WDs are not visible, due to the cooling process of WDs. Already at 5 Gyr, all WDs with masses higher than $\sim 0.55$M are now undetectable.
4.1.1 Initial binary fraction of 10%
@@
We start by looking at the 10% initial binary fraction plots in Fig. 2. At 2 Gyr the separations from all MOCCA datasets are slightly too wide and the masses too high. At 5 Gyr, escapers and isolated binaries agree very well with the observations in separation. The in-cluster binaries at this time are biased toward smaller separations than the other datasets. At 9 Gyr, not much changes compared to 5 Gyr: the maximum mass is slightly lower because more binaries are nondetectable due to increasing magnitude as they cool down. However, for all snapshots, the masses are not well reproduced. We are not able to produce very low mass WDs with mass below 0.2 M in our simulations. Instead, our close WDs usually have a mass between 0.25 and 0.4 M. There are a few observed WDs in this mass range as well, but the large majority are lower mass.
@@
These ELMs cannot form through common envelope evolution, which produces a problem for our comparisons. Currently, BSE does not handle this kind of evolution well. There is a large uncertainty in the magnetic braking, mass stability criteria, and mass transfer modeling. More precisely, to form these objects we need dynamically stable mass transfer (e.g., Belloni & Schreiber 2023b). During this stable mass transfer involving a Roche-lobe filling subgiant or an unevolved red giant star, due to strong magnetic braking, the evolution is convergent and the binary evolves toward shorter orbital periods. This kind of mass transfer is not currently handled by BSE or any other population synthesis code. In order to be able to properly reproduce the distribution of mass for these kinds of ELMs, upgrades to our BSE are needed; however, this is outside the scope of this paper.
@@
Escaping binaries are able to form lower mass WDs than in-cluster or isolated binaries, which can be seen as a subpopulation between 0.3 and 0.45 M. This is most likely due to dynamical interactions before they escaped their host cluster. Either the in-cluster binaries that would form these low mass WDs are escaping or dynamical interactions are causing an exchange, a break-up, or a merger.
@@
At $\sim 0.51$ $\rm M_{\odot}$ we can see a soft limit for WD masses. There are WDs with masses lower than this, but not many. This is a limit in BSE and a lower mass would require processes that are neither in BSE nor solved in theory. This includes, as mentioned above, uncertainty in magnetic braking, mass stability criteria, and mass transfer modeling.
4.1.2 Initial binary fraction of 95%
@@
For a 95% initial binary fraction (Fig. 3) we have a similar scenario, but there are small differences. At 2 Gyr we see a larger shift toward higher separations compared to a 10% binary fraction. As expected, similarly to the 10% binary fraction, we are not able to form the ELMs WDs that have been observed. At 5 and 9 Gyr, we have good agreement in separation between escapers, isolated binaries, and observations. In-cluster binaries are biased toward lower separations.
@@
Looking at the number of close binaries we can see that the number of binaries is increasing with time for escapers and isolated binaries since over time more binaries escape their host cluster or the components of the binary evolve into WDs. However, for in-cluster binaries, the number is decreasing due to escaping binaries and dynamical interactions causing break-ups, mergers, or exchanges.
@@
An interesting feature is a triangular gap in Fig. 3, panel (a). This gap occurs between 0.02 and 0.07 AU in projected separation and 0.8 and 1.1 $\rm M_{\odot}$ in mass. This gap is present for all datasets and the reason for it is currently not completely understood. However, it is most likely due to the pairing of objects in binaries when the initial cluster model is set up (Kiminki & Kobulnicky 2012; Sana et al. 2012; Kobulnicky et al. 2014).
Refer to captionRefer to captionRefer to caption
@@
Figure 2: Projected separation plotted against mass of the lower mass WD in the binary for close binaries inside clusters with 10% initial binary fraction. See Fig. 1 for an explanation of panels, markers, and colors.
Refer to captionRefer to captionRefer to caption
@@
Figure 3: Projected separation plotted against mass of the lower mass WD in the binary for close binaries inside clusters with a 95% initial binary fraction. See Fig. 1 for an explanation of panels, colors, and markers